The radioactive nuclei 60Co is widely used in medical applications. It undergoes
beta decay, and the total energy of the decay process is 2.82 MeV per decay event. The
half-life of this nucleus is 272 days. Suppose that a patient is given a dose of 6.9 µCi of
60Co. If all of this material decayed while in the patient's body, what would be the total
energy deposited there?
1
Expert's answer
2020-01-03T09:22:10-0500

"1Ci=" "3.7*10^{10} deacys\/sec"

Number of decays per second in 6.9"\mu" Ci of cobalt-60"=6.9*10^{-6}*3.7*10^{10}=25.53*10^4"

As, it takes approximately 5 half life for the complete decaying of a radioactive nucleus

So total number of decays in time span of 5 half life "=5*272*24*60*60*25.53*10^4=3*10^{13}"

energy released per decay event is 2.72 MeV or 2.72*106 eV

so, energy released in time span of 5 half life =3*1013 *2.72*106 eV= 8.16*1019 eV

The total energy deposited is 8.16*1019 eV